Provide symbols (where appropriate) and definitions for the following terms:
(a) Enthalpy change
(b) Entropy change
(c) Free-energy change
(d) Spontaneity process
(e) Non-spontaneous process
1
Expert's answer
2012-10-10T08:42:35-0400
(a) An enthalpy change, ΔH, describes the change in enthalpy observed in the
constituents of a thermodynamic system when undergoing a transformation or
chemical reaction. It is the difference between the enthalpy after the process
has completed, i.e. the enthalpy of the products, and the initial enthalpy of
the system, i.e. the reactants.

(b) The entropy change of a system may
be defined as the integral of all the terms involving heat exchanged (q) divided
by the absolute temperature (T) during each infinitesimally small change of the
process carried out reversibly at constant temperature.
ΔS = Sfinal –
Sinitial = qrev/T
If heat is absorbed, then ΔS = +ve and if heat is evolved,
then ΔS = –ve

(c) The free energy change of a chemical process under
standard state conditions, ΔG, can be determined four different ways:
1. If
we know the standard free energy changes of formation, ΔG(f), of each species in
a change we can determine the standard state free energy change, ΔG, for the
change using the following equation:
ΔG = ΣΔG(products) -
ΣΔG(reactants)
2. If we know the enthalpy change, ΔH, and the entropy change,
ΔS, for a chemical process, we can determine the standard state free energy
change, ΔG, for the process using the following equation:
ΔG = ΔH - TΔS
3.
If we know the equilibrium constant, Keq, for a chemical change (or if we can
determine the equilibrium constant), we can calculate the standard state free
energy change, Go, for the reaction using the equation:
ΔG = - R * T *
lnKeq
4. If we know the standard state cell potential, Eo, for an
electrochemical cell (or if we can determine the standard state cell potential),
we can calculate the standard state free energy change, Go, for the cell
reaction using the equation:
ΔG = - n * F * Eo
(d) A spontaneous process
is a process that occurs in a system without any input of energy from the
surroundings.
(e) A nonspontaneous process is an unnatural process that
after it is started will not continue to happen without outside intervention. A
nonspontaneous process only happens when outside action introduces energy to
drive the process.